Part A


Use the information below to answer the questions that follow: You have been granted a loan of R50 000 payable monthly over a 5-year term. The interest rate charged is 24% per annum compounded monthly. You are required to determine:

i. the monthly payment for this loan. [1]

ii. the total interest that will be paid after 5 years of making monthly payments [2]

iii. the monthly payment if the loan has a residual of 20% [2]

iv. the total interest paid after 5 years of making payments that consider the residual [2]

v. (ignoring the residual) what is the initial payment of the loan if you are willing to make a growing contribution of R50 every month starting from month 2? [3]



1
Expert's answer
2022-04-06T15:49:04-0400

12*5 = 60months

5000 = 1 month

x= 60 months

x = 60*5000 = 300000